### v2

@@ 5.3 Numerical methods for concentrated emulsions
Influence of drop coalescence and breakup Transient evolution of the emulsion micro-structure in concentrated emulsions including changes in droplet topology (e.g., breakup and coalescence events) remains an open area of research. The critical effect of flow-induced droplet breakup and fragmentation on the microstructure and rheology of emulsions [120, 154, 155], including wall-effects, external force fields, viscoelastic contributions have been reviewed or studied elsewhere [156, 154, 157, 77, 158, 159, 160].
Coalescence and breakup events may coexist in confined emulsion flows leading to non-trivial rheology. For example, shear bands which are regions of high and low droplet concentrations in the vorticity and flow direction, respectively, have been observed in numerical experiments[161, 148]. Figure 5(g) shows a snapshot of the droplet microstructure in VoF simulations adapted from Ref. [148]. Rosti et. al [149] determined the effective viscosity of concentrated emulsions Using a 3-dimensional VoF method for volume fractions in the range $0.310^{-3}<\phi<$ 0.3 and capillary number, $0.30.1<Ca<$ 0.3. Coalescence events lead to a non-monotonic variation of effective viscosity with $\phi$, with a peak around $\phi\approx 0.20$. Representative droplet shape distribution observed in their VoF simulation is shown in Fig. 5(h). Recently, Girotto et al. [146] used mesoscopic Lattice-Boltzmann method to study the evolution of the microstructure of emulsions as the disperse-phase volume fraction increases from semi-dilute to jammed configurations. The authors included coalescence and breakup events and further studied ageing dynamics effects after the flow is stopped. An evolution of the emulsion droplet network as the concentration increases is shown in Fig. 5(e). For a comprehensive review on numerical aspects and recent progress on the modelling of deformable particles in flows using the Lattice-Boltzmann method see Ref. [162]. Peterson et al. [163] proposed a generalized framework model for droplet breakup in dense emulsion flows using a population balance model coupled to droplet shape evolution.
@@ 6 Jammed dense emulsions with polygonal drops in a network of films
As the dispersed-phase volume fraction is increased beyond the highly concentrated regime of flowing emulsions discussed in section 5, the corresponding rheology is highly sensitive to the droplet positional structure, drop size, interparticle forces, and polydispersity. In this regime, an emulsion of repulsive droplets (stabilized against coalescence) transitions from an amorphous glass-like behavior for $\phi_{g}\approx 0.58$ to a jammed dense regime at $\phi\approx\phi_{RCP}$ where the microstructure is dense and randomly packed and $\phi_{RCP}\approx 0.64$. In the limit as $\phi\to 1$, the droplet are compressed into polygonal shapes and separated by thin films of continuous phase fluids that intersect at Plateau borders, and thus develop a microstructure or a castle of polyhedral shapes characteristic of dry foams [164, 4, 11, 1, 20, 39]. In this section, we focus on the structure and rheology of jammed dense emulsions where the droplets are densely packed showing a solid-like behavior under weak loading, and a fluid-like behavior beyond an effective yield stress [165, 37].
As discussed in sections 4-5, emulsions under flow show a non-Newtonian, viscoelastic behavior where elastic effects are typically imparted by the disperse-phase relaxation time in response to bulk stresses. Yield stress emulsions show a viscoplastic response to imposed bulk stresses. The Herschel-Bulkley model is often used for capture the flow behavior for a complex fluid that displays a yield stress, and flows with a power law relationship between stress and deformation rate above yield stress. The three parameter model including a power law exponent, $n$, consistency, $K$, and yield stress, $\tau_{Y}$, can be written as
@@
The viscoplastic behavior may be qualitatively defined by ratio of yield stress and an imposed characteristic stress,
@@
where $Bn$ is the Bingham number, $\tau_{Y}$ is the material effective yield stress, and $\tau_{c}=\mu_{B}U/L$ is the characteristic stress where $\mu_{B}$ is a viscous parameter, $U$ and $L$ are characteristic velocity and length scale, respectively.
@@
where $\sigma$ is interfacial tension coefficient, $a_{32}=3V/A$ is an volume-to-surface-area mean drop radius, and $\phi_{0}\approx 0.71$ is the limiting volume fraction at which the percolation of the droplet network collapses. The rheology of dense emulsions of non-coalescing droplets including typical flow curves and characteristic viscoelastic behavior described by the storage, $G^{\prime}$, and loss moduli, $G^{\prime\prime}$, subject to linear and non-linear viscoelastic flowing regimes has been well documented in reviews and papers[164, 20, 11, 159, 1, 51], where most of the works are experimental. Theory and numerical aspects of the problem remain an active area of research.
The measurement or observation of an apparent yield stress in jammed dense emulsions and suspension of particles with a relatively wide range of interaction is much easier than describing the underlying mechanism involving dynamics of dispersed drops in teh case of emulsions.[14, 166, 167, 168, 169, 170, 52] The collapse of the amorphous glass-like microstructure signals the transition to a fluid-like behavior where a classical empirical model by Princen and Kiss [37]for the yield stress is
@@
and $Y(\phi)$ is an empirical relation showing a logarithmic dependence on $\phi$ [37]. Several models are proposed as detailed in the review by Kim and Mason [11]. Figure 6 illustrates that two empirical models capture the trends observed experimentally for $\phi$ dependent increase modulus and yield stress. Details including the properties of dispersed and suspending fluid, the expression for computing the two quantities and values used for different constants are listed in the Appendix for completeness. For emulsions that display yield stress, recent experiments using gravity-based rheometry show the possibility of measuring both an extensional yield stress and the power law relation between extensional stress and strain rate using analysis of dripping, though challenges remain in quantitatively describing the underlying mechanisms for strong flows where droplet deformability probably plays a role.[79, 168, 23, 25]
Figure 6: Comparison between elastic modulus and yield stress empirical models for jammed dense emulsions. Data sets obtained from Refs. [171] and [172] for elastic modulus and from Refs. [37] and [38] for yield stress. Empirical models for elastic modulus obtained from Refs. [171] and [173], and for yield stress from [38] and [37]. Both elastic modulus and yield stress are normalized by a characteristic capillary stress $\sigma/a$. See Appendix B for details on the datasets and models used.
Denkov and coworkers [174] argued that the second term or the viscous stress contribution, $\tau_{v}(\dot{\gamma})$ for yielded emulsions can be attributed to the energy dissipation in thin films between neighboring drops sliding along each other. Their model anticipates a power law exponent $n$ = 1/2 if disjoining pressure is neglected, and explain why viscous stress and shear viscosity exhibit $Ca^{1/2}$ and $Ca^{-1/2}$ scaling, respectively for flowing emulsions. An extended version of the model suggests $n$ < 1/2 if interfacial dissipation plays a role and $n$ > 1/2 if disjoining pressure exerts an influence. The model appears to capture the diversity in power law exponents observed experimentally in flowing emulsions [175, 174].
Numerical studies of jammed dense emulsions Emulsions display $\phi$ dependent yield stress, and is often used by experimentalists as a model system for investigating rheological response. Numerically modeling jammed dense emulsions proffers a similar opportunity with the advantage that changes in microstructure below and above yield stress in response to applied stress can be visualized and analyzed, as shown in a recent numerical investigation by Negro et al. [147]. The authors numerically investigated in 2D the yield stress and flow behavior of a model emulsion that contains an amorphous deformable, non-coalescing droplets embedded in a Newtonian fluid, as summarized below.
@@ 7 Challenges, opportunities, and prognosis
Over the past century, the progress in describing the physicochemical origins of the flow behavior of emulsions reflects progress in describing soft matter physics, thermodynamics, intermolecular and surface forces, interfacial properties, and drop deformation, breakup and coalescence. Despite the progress, designing more sustainable, cost-efficient, or functional formulations in form of emulsions remains challenging as many fundamental scientific problems arise. The macromcolecular, supramolecular and particulate ingredients can alter the rheology of dispersed or suspending fluids and influence interfacial properties, affecting stability, application and processing of emulsions. The review captures some highlights from the current state-of-the-art in modeling shear rheology of emulsions containing Newtonian drops in Newtonian continuum phase with a Newtonian interface. Making any of the three non-Newtonian introduces conceptual, characterization and modeling challenges. Additional open questions are encountered in the following contexts, where we restrict discussion to theoretical and computational challenges only.
Extensional rheology response requires carefully considering the impact of large changes in drop shape, leading to the possibility of droplet break-up and coalescence and changes in microstructure that can influence flow response [186]. For nondilute emulsions, there is also a pronounced lack of experimental data that can be used for benchmarking theoretical methods. This is partially due to challenging extensional rheology characterization with a lack of techniques to measure extensional viscosity and visualize drop shape and microstructure evolution in response to practically-relevant deformation rates [25, 23, 24].
@@
Viscoelastic suspending fluid The constitutive models and numerical studies described in this contribution are inadequate for capturing the rheological response of emulsions containing viscoelastic suspending fluids or viscoelastic droplets in a Newtonian suspending medium. The two-way coupling of bulk elastic stresses to the interfacial stress jump across the interface can be highly-nonlinear introducing challenges in modelling multiphase flows containing moving boundaries. The effect of flow-induced cross-stream migration and deformation of droplets or capsules in viscoelastic background fluids on the rheology of dilute to concentrated suspensions remains an open area of research [187, 188, 1, 77].
@@
Bubbly fluids Theoretical and experimental investigations on the transient rheology of bubble suspensions remain an active area of research [189, 190]. In the limit of emulsions containing bubbles as the dispersed phase ($\lambda\to 0$), Rust & Manga [191, 192] compared small and large deformation theories and numerical calculations to experimental results on the shape, deformation, and effective viscosity of surfactant-free bubbly suspensions.
@@
Role of deformation and processing history, including emulsification Changes in drop sizes and size distributions, and microstructure have a direct impact on the overall flow properties of emulsions. Modeling different emulsification methods and carrying out modeling with polydisperse droplets that emulate the formulations used in personal care, food, or industry-grade emulsions requires deeper dive into rheology and thermodynamics of multi-component systems [193, 81, 44, 1, 16, 3, 2, 41].
@@
Yielding and microstructural evolution of jammed dense emulsions Further research is needed to elucidate the effects of changes in local drop size, shape and number density, and topological changes involving interconnected thin films on the bulk rheology of emulsions, especially if disjoining pressure, interfacial rheology and either dispersed or suspending fluids are viscoelastic [168, 79, 25, 167, 4, 86, 52, 39, 11].
@@
We offer this survey of theoretical and numerical modeling of emulsions rheology to the scientific community, with an awareness despite this remarkable progress, many practical problems remain in producing, storing, processing, and designing emulsions. We anticipate that advances in numerical and computational methods, and emergence of exciting problems and consumer/industry driven quests involving food and personal care emulsions made with sustainable ingredients will drive the field in the near future.
@@ Appendix A Small deformation theory: clean drops distortion tensor
where $\epsilon\ll 1$, $\mathbf{A}$ is the shape distortion tensor, $a$ is the radius of the undeformed, spherical droplet, and $r=(\mathbf{x}\cdot\mathbf{x})^{1/2}$. Solution to Eqs. (2)-(3) are obtained assuming a spherical shape by, for example, superposition of vector spherical harmonics. To leading order, shape distortions are captured in the definition of the normal vector $\mathbf{n}=\nabla S(t)/|\nabla S(t)|$ such that [74]

### v3

@@ 5.4 Numerical methods for concentrated emulsions
Influence of drop coalescence and breakup Transient evolution of the emulsion micro-structure in concentrated emulsions including changes in droplet topology (e.g., breakup and coalescence events) remains an open area of research. The critical effect of flow-induced droplet breakup and fragmentation on the microstructure and rheology of emulsions [135, 186], including wall-effects [82, 83, 85, 45, 187, 87], external force fields [81, 86, 188], non-Newtonian contributions from either the dispersed or continuous phases (e.g., viscoelastic, power-law, elastoviscoplastic) have been reviewed or studied elsewhere [45, 189, 190, 191, 188].
Coalescence and breakup events may coexist in confined emulsion flows leading to nontrivial rheology. For example, shear bands which are regions of high and low droplet concentrations in the vorticity and flow direction, respectively, have been observed in numerical experiments [192, 180]. Figure 5(g) shows a snapshot of the droplet microstructure in VoF simulations adapted from Ref. [180]. Rosti et al. [181] determined the effective viscosity of concentrated emulsions using a 3-dimensional VoF method for volume fractions in the range $0.310^{-3}<\phi<$ 0.3 and capillary number $0.30.1<Ca<$ 0.3. Coalescence events lead to a non-monotonic variation of effective viscosity with $\phi$, with a peak around $\phi\approx 0.20$. The representative droplet shape distribution observed in their VoF simulation is shown in Fig. 5(h). Recently, Girotto et al. [178] used mesoscopic Lattice-Boltzmann method to study the evolution of the microstructure of emulsions as the disperse-phase volume fraction increases from semi-dilute to jammed configurations. The authors included coalescence and breakup events and further studied aging dynamics effects after the flow is stopped. An evolution of the emulsion droplet network as the concentration increases is shown in Fig. 5(e). For a comprehensive review on numerical aspects and recent progress on the modeling of deformable particles in flows using the Lattice-Boltzmann method see Ref. [193]. Peterson et al. [194] proposed a generalized framework model for droplet breakup in dense emulsion flows using a population balance model coupled to droplet shape evolution.
@@ 6 Jammed dense emulsions with polygonal drops in a network of films
On increasing the volume of the dispersed-phase or $\phi$ beyond the highly concentrated regime of flowing emulsions discussed in section 5, the rheological response shows manifestation of a yield stress, implying flow occurs only after minimum threshold value of stress (or applied force) is exceeded. The magnitude of yield stress and beyond yield stress, the flow behavior shows high sensitivity to the positional structure, size, shape, interparticle forces, and polydispersity of droplets. In this regime, an emulsion of repulsive droplets (stabilized against coalescence) transitions from amorphous glass-like behavior for $\phi_{g}\approx 0.58$ to a jammed, dense regime at $\phi\approx\phi_{RCP}$ where the microstructure is dense and randomly packed and $\phi_{RCP}\approx 0.64$. In the limit as $\phi\to 1$, the drops get compressed into polygonal shapes. The deformed drops are separated by thin films of the continuous phase fluid, and the films that intersect at Plateau borders thus develop a microstructure or a castle of polyhedral shapes characteristic of dry foams [195, 4, 8, 1, 23, 43]. In this section, we focus on the structure and rheology of jammed dense emulsions where the droplets are densely packed showing a solid-like behavior under weak loading, and a fluid-like behavior beyond an effective yield stress [196, 41].
Dilute and nondilute flowing emulsions, as discussed in sections 4-5, exhibit a non-Newtonian rheology and viscoelastic response, and their elasticity is attributed at the drop level to the interfacial tension dependent shape relaxation time. Jammed dense emulsions show a viscoplastic response to imposed bulk stresses, such that flow only occurs after yield stress is exceeded. The empiricial Herschel-Bulkley model is often used for capturing the flow behavior for a complex fluid that displays a yield stress and flows with a power law relationship between stress and deformation rate above yield stress. The three parameter HB model includes a power law exponent, $n$, consistency, $K$, and yield stress, $\tau_{Y}$, and can be written as
@@
The HB model is a generalization of the Bingham model that includes only two parameters (as the power law exponent equals 1). More elaborate models such as SGR (soft glassy rheology) model, mode coupling theory can be derived to capture the rheological behavior of yield stress materials by accounting for local traps or local rearrangement zones in dispersions containing densely packed drops or particles, as detailed elsewhere [197, 4, 198].
The viscoplastic behavior may be qualitatively defined using the Bingham number, $Bn$
@@
which is simply the ratio of yield stress $\tau_{Y}$ and an imposed characteristic stress, $\tau_{c}=\mu U/L$. Here $\mu$ is characteristic viscosity, $U$ and $L$ are characteristic velocity and length scale, respectively.
@@
where $\sigma$ is interfacial tension coefficient, $a_{32}=3V/A$ is an volume-to-surface-area mean drop radius, and $\phi_{0}\approx 0.71$ is the limiting volume fraction at which the percolation of the droplet network collapses. The rheology of dense emulsions of non-coalescing droplets including typical flow curves and characteristic viscoelastic behavior described by the storage, $G^{\prime}$, and loss moduli, $G^{\prime\prime}$, subject to linear and non-linear viscoelastic flowing regimes and has been well documented in reviews and papers[195, 23, 8, 187, 1, 56], where most of the works are experimental. Theory and numerical aspects of the problem remain an active area of research.
The measurement or observation of an apparent yield stress in jammed dense emulsions and suspension of particles with a relatively wide range of interaction is much easier than describing the underlying mechanism involving dynamics of dispersed drops in the case of emulsions.[7, 199, 197, 200, 201, 202, 57] The collapse of the amorphous glass-like microstructure signals the transition to a fluid-like behavior where a classical empirical model by Princen and Kiss [41] for the yield stress is
@@
and $Y(\phi)$ is an empirical relation showing a logarithmic dependence on $\phi$ [41]. Several models are proposed as detailed in the review by Kim and Mason [8]. Figure 6 illustrates that two empirical models capture the trends observed experimentally for $\phi$ dependent increase in modulus and yield stress. Details including the properties of dispersed and suspending fluid, the expression for computing the two quantities and values used for different constants are listed in the Appendix for completeness. For emulsions that display yield stress, recent experiments using gravity-based rheometry show the possibility of measuring both an extensional yield stress and the power law relation between extensional stress and strain rate using analysis of dripping, though challenges remain in quantitatively describing the underlying mechanisms for strong flows where droplet deformability probably plays a role.[89, 200, 26, 28]
Figure 7: Comparison between elastic modulus and yield stress empirical models for jammed dense emulsions. Data sets obtained from Refs. [203] and [204] for elastic modulus and from Refs. [41] and [42] for yield stress. Empirical models for elastic modulus obtained from Refs. [203] and [205], and for yield stress from [42] and [41]. Both elastic modulus and yield stress are normalized by a characteristic capillary stress $\sigma/a$. See Appendix C for details on the datasets and models used.
Denkov and coworkers [206] argued that the second term or the viscous stress contribution, $\tau_{v}(\dot{\gamma})$ for yielded emulsions can be attributed to the energy dissipation in thin films between neighboring drops sliding along each other. Their model anticipates a power law exponent $n$ = 1/2 if disjoining pressure is neglected, and explains why viscous stress and shear viscosity exhibit $Ca^{1/2}$ and $Ca^{-1/2}$ scaling, respectively for flowing emulsions. An extended version of the model suggests $n$ < 1/2 if interfacial dissipation plays a role and $n$ > 1/2 if disjoining pressure exerts an influence. The model appears to capture the diversity in power law exponents observed experimentally in flowing emulsions [207, 206].
Numerical studies of jammed dense emulsions Emulsions display $\phi$ dependent yield stress, and is often used by experimentalists as a model system for investigating rheological response. Numerically modeling jammed dense emulsions proffers a similar opportunity with the advantage that changes in microstructure below and above yield stress in response to applied stress can be visualized and analyzed, as shown in a recent numerical investigation by Negro et al. [179]. The authors numerically investigated in 2D the yield stress and flow behavior of a model emulsion that contains an amorphous deformable non-coalescing droplets embedded in a Newtonian fluid, as summarized below.
@@ 7 Challenges, opportunities, and prognosis
Over the past century, the progress in describing the physicochemical origins of the flow behavior of emulsions reflects progress in describing soft matter physics, thermodynamics, intermolecular and surface forces, interfacial properties, and drop deformation, breakup and coalescence. Despite progress, designing more sustainable, cost-efficient, or functional emulsion-based formulations remains challenging as many fundamental scientific problems arise. The macromolecular, supramolecular and particulate ingredients can alter the rheology of dispersed or suspending fluids and influence interfacial properties, affecting stability, application and processing of emulsions. The review captures some highlights from the current state-of-the-art in modeling shear rheology of emulsions containing Newtonian drops in Newtonian continuum phase with a Newtonian interface. Making any of the three non-Newtonian introduces conceptual, characterization and modeling challenges. Additional open questions are encountered in the following contexts, where we restrict discussion to theoretical and computational challenges only.
Figure 8: A compendium of analytical expressions for relative viscosity for dilute to nondilute emulsions derived using small deformation or effective medium theory and phenomenological relations for yield stress in jammed dense emulsions. The publication years are included for each model to highlight the milestones corresponding to time’s arrow.
Extensional rheology response requires a careful consideration of large changes in drop shapes, which enhances the possibility of breakup or coalescence of drops and microstructure changes, that in turn influences the response of streamwise velocity gradients[218]. For nondilute emulsions, there is also a pronounced lack of experimental data that can be used to benchmark theoretical methods. There is a lack of in-situ techniques that can be used to measure extensional viscosity while visualizing the evolution of drop shapes and microstructure in response to practically relevant deformation rates [28, 26, 27].
@@
Viscoelastic dispersed or suspending fluid Despite significant progress in analyzing drop deformation for cases with either phase or both phases are non-Newtonian, the constitutive models and numerical studies described in this contribution are inadequate for capturing the rheological response at high deformation rates for emulsions containing viscoelastic suspending fluids or viscoelastic droplets in a Newtonian suspending medium. The two-way coupling of bulk elastic stresses to the interfacial stress jump across the interface can be highly nonlinear, introducing challenges in modeling multiphase flows containing moving boundaries. The effect of flow-induced cross-stream migration and deformation of droplets or capsules in viscoelastic background fluids on the rheology of dilute to concentrated suspensions also remains an open area of research [219, 220, 1, 45].
@@
Bubbly fluids Theoretical and experimental investigations on the transient rheology of bubble suspensions remain an active area of research [221, 222]. In the limit of emulsions containing bubbles as the dispersed phase ($\lambda\to 0$), Rust & Manga [223, 224] compared small and large deformation theories and numerical calculations to experimental results on the shape, deformation and effective viscosity of surfactant-free bubbly suspensions. Around the same time, Llewellin et al. [225] proposed a constitutive model to study the transient rheology of polydisperse bubble suspensions derived using the small deformation theory of Frankel & Acrivos [15] that reduces to a three-parameter Jeffrey-like model typically used to characterize the rheological response of viscoelastic liquids, as shown in Eq. (60) [10].
@@
Role of deformation and processing history, including emulsification Changes in drop sizes and size distributions, and microstructure have a direct impact on the overall flow properties of emulsions. Modeling different emulsification methods [103, 104, 33, 34, 154] and polydisperse droplets to emulate the formulations used in personal care, food, or industry grade emulsions requires a deeper dive into the rheology and thermodynamics of multicomponent systems [226, 91, 49, 1, 18, 3, 2, 46].
@@
Yielding and microstructural evolution of jammed dense emulsions Further research is needed to elucidate the effects of changes in local drop size, shape and number density, and topological changes involving interconnected thin films on the bulk rheology of emulsions, especially if disjoining pressure, and interfacial rheology effects are to be included, and if non-Newtonian phases are involved. [200, 89, 28, 197, 4, 96, 57, 43, 8].
@@
We close this overview with Figure 8 that highlights the key equations and the underlying physicochemical considerations described in this contribution. Figure 8 includes a timeline of analytical relative viscosity models from dilute to semi-dilute systems and phenomenological relations for yield stress of jammed dense emulsions. We offer this survey of theoretical and numerical modeling of emulsions rheology to the scientific community, with the awareness that despite this remarkable progress, many practical problems remain in producing, storing, processing, and designing emulsions. We anticipate that advances in numerical and computational methods, and the emergence of exciting problems and consumer/industry-driven quests involving food and personal care emulsions made with sustainable ingredients will drive the field in the near future.
Appendix
@@ Appendix A Small deformation theory: clean drops distortion tensor
where $\epsilon\ll 1$ is a small parameter, $\mathbf{A}$ is the shape distortion tensor, $a$ is the radius of the undeformed, spherical droplet, and $r=(\mathbf{x}\cdot\mathbf{x})^{1/2}$ is the radial position measured from the droplet center. The shape distortion tensor is a measure of elongations of material lines of a deformable body,
$\frac{ds}{ds_{0}}=\left(1-2\epsilon\frac{\mathbf{x}}{r}\cdot\mathbf{A}\cdot\frac{\mathbf{x}}{r}\right)^{-1/2}\,,$ (80)
where $ds_{0}$ and $ds$ are the length of a material line at a reference and current configuration. By definition, the shape distortion tensor $\mathbf{A}=\left(\mathbf{I}-\mathbf{B}^{-1}\right)/2$, where $\mathbf{B}=\mathbf{F}\cdot\mathbf{F}^{T}$ is the left Cauchy-Green deformation tensor and $\mathbf{F}$ is the deformation gradient tensor. The deformation gradient tensor is a second order tensor that maps the deformation of material lines from a reference to a current configuration $d\mathbf{x}=\mathbf{F}\cdot d\mathbf{x}_{0}$ whose components are $F_{ij}=\partial x_{i}/\partial x_{0j}$ [145]. Taking a Taylor series expansion of the right-hand-side of Eq. (80) yields,
$\frac{ds}{ds_{0}}=1+\epsilon\frac{\mathbf{x}}{r}\cdot\mathbf{A}\cdot\frac{\mathbf{x}}{r}+\epsilon^{2}\frac{3}{2}\frac{\mathbf{x}\mathbf{x}:\mathbf{A}\mathbf{A}:\mathbf{x}\mathbf{x}}{r^{4}}+\ldots\,.$ (81)
Equation (79) is obtained from relation (81) noting that perturbations in the droplet shape are captured by $r(t)=a(ds/ds_{0})$ and hence the surface is defined by $S(t)=r(t)-ds/ds_{0}\equiv 0$ to leading order in $\epsilon$.
The solutions to Eqs. (2)-(3) are obtained assuming a spherical shape by, for example, superposition of vector spherical harmonics [102]. To leading order, shape distortions are captured in the definition of the normal vector $\mathbf{n}=\nabla S(t)/|\nabla S(t)|$ such that [78]